# Content Accuracy and Quality Aware Resource Allocation Based on LP-Guided DRL for ISAC-Driven AIGC Networks

### v1

@@ Abstract
Integrated sensing and communication (ISAC) can enhance artificial intelligence-generated content (AIGC) networks by providing efficient sensing and transmission. Existing AIGC services usually assume that the accuracy of the generated content can be ensured, given accurate input data (e.g., pose image) and command (i.e., prompt), thus only the content generation quality (CGQ) is concerned. However, it is not applicable in ISAC-based AIGC networks, where content generation is based on inaccurate sensed data. Moreover, the AIGC model itself introduces generation errors, which depend on the number of generating steps (i.e., computing resources). Thus, to assess the quality of experience (QoE) of ISAC-based AIGC services, this paper proposes a content accuracy and quality aware service assessment metric (CAQA). Since allocating more resources to sensing and generating improves content accuracy but may reduce communication quality, and vice versa, this sensing-generating (computing)-communication three-dimensional resource tradeoff must be optimized to maximize the average CAQA (AvgCAQA) across all users with AIGC (CAQA-AIGC). This problem is NP-hard, with a large solution space that grows exponentially with the number of users. To solve the CAQA-AIGC problem with low complexity, a standard linear programming (LP) guided deep reinforcement learning (DRL) algorithm with an action filter (LPDRL-F) is proposed. Through the LP-guided approach and the action filter, LPDRL-F can transform the original three-dimensional solution space to two dimensions, reducing complexity while improving the learning performance of DRL. Simulations show that compared to existing DRL and generative diffusion model (GDM) algorithms without LP, LPDRL-F converges faster by over 60% and finds better resource allocation solutions, thus improving AvgCAQA by more than 14%. With LPDRL-F, CAQA-AIGC can achieve an improvement in AvgCAQA of more than 50% compared to existing schemes focusing solely on CGQ.
Index Terms:
ISAC, AIGC service quality, generation error, content accuracy and quality aware, sensing-computing-communication, LP-guided deep reinforcement learning.
1 Introduction
@@
  • We conduct extensive simulations to verify the effectiveness of the proposed schemes. It will be shown that, compared to existing DRL and generative diffusion model (GDM) algorithms, LPDRL-F converges faster by over 60% and finds better resource allocation solutions, thus improving AvgCAQA by more than 14%. Moreover, compared to the CGQ-only oriented schemes, the proposed CAQA-AIGC with LPDRL-F improves AvgCAQA by over 50%.
The remainder of the paper is organized as follows. Section 2 introduces the system model. The CAQA-AIGC problem of maximizing AvgCAQA is formulated in Section 3. Section 4 details the LPDRL-F algorithm. Section 5 presents simulation results. Finally, conclusions are drawn in Section 6.
2 System Model
2.1 ISAC-Based AIGC Network Architecture
@@
2.2 ISAC-Based AIGC Service Workflow and Performance Modeling
@@
3 CAQA-AIGC Problem Formulation
3.1 QoE Model for ISAC-Based AIGC Service
@@ 3.2 Problem Formulation
The CAQA-AIGC problem is non-convex and NP-hard due to the nonlinear relationship between sensing energy and sensing accuracy, as well as the dependency of AEG on the number of generation steps. Moreover, the resource coupling caused by queuing times for computing during the generation phase (Phase II) and the waiting times for communication before the communication phase (Phase III) jointly affect the service time. This intrinsic coupling significantly increases the complexity of the overall resource allocation problem. Note that given the sensing energy and generating step solutions, the allocation of communication energy can be transformed into a standard LP problem. However, sensing and generating resource allocation (SGenRA) remains a non-convex and NP-hard problem. Considering these properties, combined with the fact that DRL can deal with nonconvex problems [27], a low complexity LP-guided DRL-F (LPDRL-F) algorithm is proposed to solve the CAQA-AIGC problem, which reduces the solution space through the LP-guided approach and the action filter.
4 LP-Guided DRL-F Algorithm
4.1 Structure and Main Stages
@@
The structure of the proposed LPDRL-F is shown in Fig. 4. Note that the joint optimization problem is divided into two sub-problems, i.e., SGenRA, and ComRA. First, a DRL-F algorithm is proposed to allocate sensing energy and generating steps based on user states, including channel condition $h_{k}$} and CAQA requirement $\Omega_{k,\min}$}. Given the sensing energy ${\bf{E}}_{s}$} and generating step $\bf{z}$} allocation solutions, the generating accuracy $\Theta_{k}$} can be obtained. Moreover, ComRA sub-problem can be analytically demonstrated as a standard LP problem. Then, users are ranked based on their generation accuracy and channel conditions, and a ranking-based communication energy allocation algorithm (RCE) is proposed to obtain the optimal communication energy allocation solution. Given the allocated communication energy, the communication capacity can be calculated. Finally, AvgCAQA is obtained based on the sensing energy and generating step solutions from DRL-F and the communication energy solution from RCE.
@@
Note that in LPDRL-F, the deep neural network (DNN), including shared hidden layers and two separate branches, i.e., the sensing branch and the step branch, should be trained to learn the policy mapping from user states to sensing energy and generating step actions. To enhance the generalization capability of the proposed algorithm, user states with a total size of 2KK, including $K$ channel conditions and $K$ CAQA requirements, are dynamically changing at each iteration. Specifically, the CAQA requirements follow a uniform distribution [17]. At each iteration, user states are input to shared hidden layers of the DNN to extract a shared feature vector $s_{i,s}$}, which is then fed to two separate branches. Since the generating steps are discrete, Gumbel-softmax method [28] is employed to obtain the generating step solutions for $K$ users, i.e., ${\bf{z}}(i)$. These discrete steps are concatenated with the shared features to form the input of the sensing branch. This design enables the sensing branch to capture the feature of the generating steps, thereby ensuring that the output original sensing energy actions are well adapted to the action filter determined by the generating steps. The sensing branch employs a multi-layer perceptron (MLP) with sigmoid activation [29] to obtain $K$ continuous values, representing the original sensing energy actions, which are processed by the action filter to obtain sensing energy solutions, i.e., ${\bf{E}}_{s}$}. Given joint sensing and generating solutions, i.e., $\left\{{{{\bf{E}}_{s}}\left(i\right),{\bf{z}}\left(i\right)}\right\}$, CAQA-AIGC problem can be reduced to ComRA sub-problem. This is an LP problem, which can be effectively solved by the RCE algorithm with low complexity. Thus, the communication energy allocation solution ${\bf{E}}_{c}$} can be obtained, and the value of AvgCAQA, i.e., $\frac{1}{K}\sum\nolimits_{k=1}^{K}{{\eta_{k}}\left({{E_{k,s}},{E_{k,c}}}\right)}$ can be calculated and fed back to the DNN. This feedback, user states, and joint sensing energy and generating step action are used as training data to guide the DNN to refine its policy, enabling it to maximize AvgCAQA and optimize the joint sensing, generating, and communication resource allocation strategy over time. Therefore, LPDRL-F involves three stages:
@@
(1) Initialization Phase (Blue and red arrow flow): The DNN in LPDRL-F is initialized via orthogonal initialization [30], which can effectively mitigate the problem of vanishing or exploding gradients. Note that action filter does not require initialization because it is rule-based (see Section 4.2) and determined by the user’s generating steps and CAQA requirements. Similarly, the RCE algorithm also does not require initialization. At the beginning of the algorithm, the replay buffer $\Psi$ is initialized to be empty. Given the observed states $s_{i}$} of all users in the $i$-th iteration, including the channel condition $h_{k}(i)$ and CAQA requirements ${\Omega_{k,\min}}\left(i\right)$, the joint action $a_{i}$}, including the generating step ${\bf{z}}\left(i\right)$ and original sensing energy allocation action, is generated by the DNN. The sensing energy allocation solutions, i.e., ${{\bf{E}}_{s}}\left(i\right)$, is obtained by the action filter. Then, given ${\bf{z}}\left(i\right)$ and ${{\bf{E}}_{s}}\left(i\right)$, the communication energy allocation solution ${{\bf{E}}_{c}}\left(i\right)$ can be obtained by the RCE algorithm. According to (9), AvgCAQA is obtained with ${\bf{z}}\left(i\right)$, ${{\bf{E}}_{s}}\left(i\right)$, and ${{\bf{E}}_{c}}\left(i\right)$, it is taken as a reward $r_{i}$} for DNN. Then, $\left({{s_{i}},{a_{i}},{r_{i}},{s_{i+1}}}\right)$ (see Section 4.2 for detailed definitions), is input to the replay buffer $\Psi$ as a single experience data. This process continues until the replay buffer accumulates sufficient data to commence training, such as ${\Psi_{\min}}=5000$ experiences.
@@
To further illustrate the proposed algorithm, the detailed designs of the DRL-F algorithm for SGenRA and the RCE algorithm for ComRA are presented in the following subsections.
4.2 DRL-F for SGenRA Sub-problem
@@
(3) LP-guided reward: Given the generating step and sensing energy allocation solutions from the DRL-F algorithm, the communication energy allocation ${\bf{E}}_{c}$} is determined by the RCE algorithm, as outlined in Section 4.3. AvgCAQA is computed based on sensing energy, generating step, and communication energy allocations. Since constraints C3 and C4 are satisfied in action design, and constraints C1, C2, and C6 can be satisfied in the RCE algorithm, the LP-guided reward $r$ is defined to ensure compliance with constraint C5. Specifically, if all users satisfy the individualized minimum CAQA requirement, i.e., ${\Omega_{k}}\geq{\Omega_{k,\min}}$, the reward is the AvgCAQA. Otherwise, a penalty is introduced based on the number of users violating C5. The reward function is given by
@@
A specific DRL algorithm, i.e., the soft actor-critic (SAC) framework, is employed in DRL-F due to its capability to provide improved training stability through entropy regularization [32, 33], which consists of two critic networks, $Q_{{\theta_{1}}}$} and $Q_{{\theta_{2}}}$}, and an actor network $\pi_{\phi}$}. At iteration $i$, the algorithm aims to maximize the weighted cumulative discounted reward and strategy entropy [32], as follows:
@@
where $\left|{{\Psi_{b}}}\right|$ is the number of batch and $y_{i}$} is the target value of critic network, given by ${y_{i}}={r_{i}}+\gamma\left({\mathop{\min}\limits_{j=1,2}\left({{Q_{{{\theta^{\prime}}_{j}}}}({s_{i+1}},{a_{i+1}})-\rho{\rm{lo}}{{\rm{g}}_{{\pi_{\phi}}}}({{\tilde{a}}_{i+1}}|{s_{i+1}})}\right)}\right)$, where ${\tilde{a}_{i+1}}={\pi_{\phi}}\left({\cdot|{s_{i+1}}}\right)$ emphasizes that the next action should be resampled from the policy.
The temperature parameter $\rho$ reveals the relative importance of reward versus entropy term, brings the randomness of the optimal policy, and can be adaptively optimized according to ${\nabla_{\rho}}J\left(\rho\right)$ [29]. The actor network parameters are updated according to [32]
@@
4.3 RCE Algorithm for ComRA Sub-problem
@@
where constraints M1 is derived from the original C1, C5, and C6 constraints, and M2 is derived from the original C2 constraint in the CAQA-AIGC problem.
@@
ComRA sub-problem aims to optimize the communication energy allocation for each user to maximize the content generation quality (CGQ). To solve this problem, the RCE algorithm is proposed, as detailed in Algorithm 1 (see Appendix A). The optimality of the RCE algorithm is established in the following theorem.
@@
The proof of Theorem 1 is provided in Appendix A.
4.4 Complexity Analysis
For ComRA sub-problem, the computational complexity of standard LP solvers, such as interior-point [34] and simplex [35] methods, is theoretically bounded by $O\left({{K^{3.5}}}\right)$ (polynomial) and $O\left({{2^{K}}}\right)$ (exponential), respectively. In contrast, the proposed RCE algorithm exploits the inherent structure of ComRA sub-problem, i.e., the linear objective and separable box constraints, to achieve a complexity of $O\left({K\log K}\right)$ (see Appendix A). Notably, the complexity reduction does not sacrifice optimality. The allocation strategy provably converges to the global optimum with linear objectives and box constraints (see Theorem 1). This efficiency enables real-time energy allocation in ISAC, where traditional LP tools (e.g., CVXPY, MATLAB’s linprog) become intractable, especially for larger $K$.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) can enhance artificial intelligence-generated content (AIGC) networks by providing efficient sensing and transmission. Existing AIGC services usually assume that the accuracy of the generated content can be ensured, given accurate input data (e.g., pose image) and command (i.e., prompt), thus only the content generation quality (CGQ) is concerned. However, it is not applicable in ISAC-based AIGC networks, where content generation is based on inaccurate sensed data. Moreover, the AIGC model itself introduces generation errors, which depend on the number of generating steps (i.e., computing resources). Thus, to assess the quality of experience (QoE) of ISAC-based AIGC services, this paper proposes a content accuracy and quality aware service assessment metric (CAQA). Since allocating more resources to sensing and generating improves content accuracy but may reduce communication quality, and vice versa, this sensing-generating (computing)-communication three-dimensional resource tradeoff must be optimized to maximize the average CAQA (AvgCAQA) across all users with AIGC (CAQA-AIGC). This problem is NP-hard, with a large solution space that grows exponentially with the number of users. To solve the CAQA-AIGC problem with low complexity, a standard linear programming (LP) guided deep reinforcement learning (DRL) algorithm with an action filter (LPDRL-F) is proposed. Through the LP-guided approach and the action filter, LPDRL-F can transform the original three-dimensional solution space to two dimensions, reducing complexity while improving the learning performance of DRL. Simulations show that compared to existing DRL and generative diffusion model (GDM) algorithms without LP, LPDRL-F converges faster and finds better resource allocation solutions, thus improving AvgCAQA by more than 10%. With LPDRL-F, CAQA-AIGC can achieve an improvement in AvgCAQA of more than 50% compared to existing schemes focusing solely on CGQ.
publicationid: pubid: 1536–1233 © 2026 IEEE. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information.
I Introduction
@@
  • We conduct extensive simulations to verify the effectiveness of the proposed schemes. It will be shown that, compared to existing DRL and generative diffusion model (GDM) algorithms, LPDRL-F converges faster and finds better resource allocation solutions, thus improving AvgCAQA by more than 10%. Moreover, compared to the CGQ-only oriented schemes, the proposed CAQA-AIGC with LPDRL-F improves AvgCAQA by over 50%.
The remainder of the paper is organized as follows. Section II introduces the system model. The CAQA-AIGC problem of maximizing AvgCAQA is formulated in Section III. Section IV details the LPDRL-F algorithm. Section V presents simulation results. Finally, conclusions are drawn in Section VI.
II System Model
II-A ISAC-Based AIGC Network Architecture
@@
II-B ISAC-Based AIGC Service Workflow and Performance Modeling
@@
III CAQA-AIGC Problem Formulation
III-A QoE Model for ISAC-Based AIGC Service
@@
In the CAQA model, content accuracy reflects sensing accuracy and generation errors, which directly impact user-perceived quality because they determine whether the generated result matches the desired one. It primarily influences the visual fidelity of regions of interest (ROI) and desired consistency [27, 28], as users are sensitive to whether the desired consistency of ROI regions (e.g., pose consistency) adheres to their intended expectations. Content quality is related to the delivered resolution, which shapes the visual fidelity and fine-grained structural details within ROI regions. Accordingly, content quality mainly corresponds to the psychological factors of visual fidelity of ROI and fine-grained details [27, 28]. Therefore, a higher CAQA can stand for a better user-perceived quality.
@@ III-B Problem Formulation
The CAQA-AIGC problem is non-convex and NP-hard due to the nonlinear relationship between sensing energy and sensing accuracy, as well as the dependency of AEG on the number of generation steps. Moreover, the resource coupling caused by queuing times for computing during the generation phase (Phase II) and the waiting times for communication before the communication phase (Phase III) jointly affect the service time. This intrinsic coupling significantly increases the complexity of the overall resource allocation problem. Note that given the sensing energy and generating step solutions, the communication energy resource allocation (ComRA) can be transformed into a standard LP problem. However, sensing and generating resource allocation (SGenRA) remains a non-convex and NP-hard problem, which can be well-solved by DRL [29]. One possible solution is to decompose the three-dimensional CAQA-AIGC problem into two sub-problems, i.e., the two-dimensional SGenRA sub-problem and the convex ComRA sub-problem. First, SGenRA can be solved by DRL. Then, based on its solution, ComRA can be solved by RCE. Thus, a low complexity LP-guided DRL-F (LPDRL-F) algorithm can be proposed to solve the CAQA-AIGC problem, which reduces the solution space through the LP-guided approach.
IV LP-Guided DRL-F Algorithm
IV-A Structure and Main Stages
@@
The structure of the proposed LPDRL-F is shown in Fig. 4. Note that the joint optimization problem is divided into two sub-problems, i.e., SGenRA, and ComRA. First, a DRL-F algorithm is proposed to allocate sensing energy and generating steps based on user states, including channel condition $h_{k}$} and CAQA requirement $\Omega_{k,\min}$}. Then, with the obtained sensing energy ${\bf{E}}_{s}$} and generating step $\bf{z}$}, the generating accuracy $\Theta_{k}$} can be obtained. Moreover, ComRA sub-problem can be analytically demonstrated as a standard LP problem, and be optimally solved by a ranking-based communication energy allocation algorithm (RCE). Thus, AvgCAQA can be obtained (see (8) and (9)).
@@
Note that in LPDRL-F, the deep neural network (DNN), including shared hidden layers and two separate branches, i.e., the sensing branch and the step branch, should be trained to learn the policy mapping from user states to sensing energy and generating step actions. To enhance the generalization capability of the proposed algorithm, user states with a total size of 2KK, including $K$ channel conditions and $K$ CAQA requirements, are dynamically changing at each iteration. Specifically, the CAQA requirements follow a uniform distribution [17]. At each iteration, user states are input to shared hidden layers of the DNN to extract a shared feature vector $s_{i,s}$}, which is then fed to two separate branches. Since the generating steps are discrete, Gumbel-softmax method [30] is employed to obtain the generating step solutions for $K$ users, i.e., ${\bf{z}}(i)$. These discrete steps are concatenated with the shared features to form the input of the sensing branch. This design enables the sensing branch to capture the feature of the generating steps, thereby ensuring that the output original sensing energy actions are well adapted to the action filter determined by the generating steps. The sensing branch employs a multi-layer perceptron (MLP) with sigmoid activation [31] to obtain $K$ continuous values, representing the original sensing energy actions, which are processed by the action filter to obtain sensing energy solutions, i.e., ${\bf{E}}_{s}$}. Given joint sensing and generating solutions, i.e., $\left\{{{{\bf{E}}_{s}}\left(i\right),{\bf{z}}\left(i\right)}\right\}$, CAQA-AIGC problem can be reduced to ComRA sub-problem, which can be effectively solved by the RCE algorithm with low complexity. Thus, the communication energy allocation solution ${\bf{E}}_{c}$} can be obtained, and the value of AvgCAQA, i.e., $\frac{1}{K}\sum\nolimits_{k=1}^{K}{{\Omega_{k}}\left({{E_{k,s}},{z_{k}},{E_{k,c}}}\right)}$ can be calculated and fed back to the DNN. This feedback, user states, and joint sensing energy and generating step action are used as training data to guide the DNN to refine its policy, enabling it to maximize AvgCAQA and optimize the joint sensing, generating, and communication resource allocation strategy over time. Therefore, LPDRL-F involves three stages:
@@
(1) Initialization Phase (Blue and red arrow flow): The DNN in LPDRL-F is initialized via orthogonal initialization [32], which can effectively mitigate the problem of vanishing or exploding gradients. Note that action filter does not require initialization because it is rule-based (see Section IV-B) and determined by the user’s generating steps and CAQA requirements. Similarly, the RCE algorithm also does not require initialization. At the beginning of the algorithm, the replay buffer $\Psi$ is initialized to be empty. Given the observed states $s_{i}$} of all users in the $i$-th iteration, including the channel condition $h_{k}(i)$ and CAQA requirements ${\Omega_{k,\min}}\left(i\right)$, the joint action $a_{i}$}, including the generating step ${\bf{z}}\left(i\right)$ and original sensing energy allocation action, is generated by the DNN. The sensing energy allocation solutions, i.e., ${{\bf{E}}_{s}}\left(i\right)$, is obtained by the action filter. Then, given ${\bf{z}}\left(i\right)$ and ${{\bf{E}}_{s}}\left(i\right)$, the communication energy allocation solution ${{\bf{E}}_{c}}\left(i\right)$ can be obtained by the RCE algorithm. According to (9), AvgCAQA is obtained with ${\bf{z}}\left(i\right)$, ${{\bf{E}}_{s}}\left(i\right)$, and ${{\bf{E}}_{c}}\left(i\right)$, it is taken as a reward $r_{i}$} for DNN. Then, $\left({{s_{i}},{a_{i}},{r_{i}},{s_{i+1}}}\right)$ (see Section IV-B for detailed definitions), is input to the replay buffer $\Psi$ as a single experience data. This process continues until the replay buffer accumulates sufficient data to commence training, such as ${\Psi_{\min}}=5000$ experiences.
@@
The pseudocode of LPDRL-F is summarized in Algorithm 1 (see Appendix A). To further illustrate the proposed algorithm, the detailed designs of the DRL-F algorithm for SGenRA and the RCE algorithm for ComRA are presented in the following subsections.
IV-B DRL-F for SGenRA Sub-problem
@@
(3) LP-guided reward: Given the generating step and sensing energy allocation solutions from the DRL-F algorithm, the communication energy allocation ${\bf{E}}_{c}$} is determined by the RCE algorithm, as outlined in Section IV-C. AvgCAQA is computed based on sensing energy, generating step, and communication energy allocations. Since constraints C3 and C4 are satisfied in action design, and constraints C1, C2, and C6 can be satisfied in the RCE algorithm, the LP-guided reward $r$ is defined to ensure compliance with constraint C5. Specifically, if all users satisfy the individualized minimum CAQA requirement, i.e., ${\Omega_{k}}\geq{\Omega_{k,\min}}$, the reward is the AvgCAQA. Otherwise, a penalty is introduced based on the number of users violating C5. The reward function is given by
@@
A specific DRL algorithm, i.e., the soft actor-critic (SAC) framework, is employed in DRL-F due to its capability to provide improved training stability through entropy regularization [34, 35], which consists of two critic networks, $Q_{{\theta_{1}}}$} and $Q_{{\theta_{2}}}$}, and an actor network $\pi_{\phi}$}. At iteration $i$, the algorithm aims to maximize the weighted cumulative discounted reward and policy entropy [34], as follows:
@@
where $\left|{{\Psi_{b}}}\right|$ represents the batch size and $y_{i}$} denotes the target value of critic network, which is given by ${y_{i}}={r_{i}}+\gamma\left({\mathop{\min}\limits_{j=1,2}\left({{Q_{{{\theta^{\prime}_{j}}}}}({s_{i+1}},{a_{i+1}})-\rho{\rm{lo}}{{\rm{g}}_{{\pi_{\phi}}}}({{\tilde{a}}_{i+1}}|{s_{i+1}})}\right)}\right)$, where ${\tilde{a}_{i+1}}={\pi_{\phi}}\left({\cdot|{s_{i+1}}}\right)$ emphasizes that the next action should be resampled from the policy.
The temperature parameter $\rho$ reveals the relative importance of reward versus entropy term, brings randomness to the optimal policy, and can be adaptively optimized according to ${\nabla_{\rho}}J\left(\rho\right)$ [31]. The actor network parameters are updated according to [34]
@@
IV-C RCE Algorithm for ComRA Sub-problem
@@
where constraint M1 is derived from the original C1, C5, and C6 constraints, and constraint M2 is derived from the original C2 constraint in the CAQA-AIGC problem.
@@
ComRA sub-problem aims to optimize the communication energy allocation for each user to maximize the content generation quality (CGQ). To solve this problem, the RCE algorithm is proposed, as detailed in Algorithm 2 (see Appendix B). The optimality of the RCE algorithm is established in the following theorem.
@@
The proof of Theorem 1 is provided in Appendix B.
IV-D Complexity Analysis
For ComRA sub-problem, the computational complexity of standard LP solvers, such as interior-point [36] and simplex [37] methods, is theoretically bounded by $O\left({{K^{3.5}}}\right)$ (polynomial) and $O\left({{2^{K}}}\right)$ (exponential), respectively. In contrast, the proposed RCE algorithm exploits the inherent structure of ComRA sub-problem, i.e., the linear objective and separable box constraints, to achieve a complexity of $O\left({K\log K}\right)$ (see Appendix B). Notably, the complexity reduction does not sacrifice optimality. The allocation strategy provably converges to the global optimum with linear objectives and box constraints (see Theorem 1). This efficiency enables real-time energy allocation in ISAC, where traditional LP tools (e.g., CVXPY, MATLAB’s linprog) become intractable, especially for larger $K$.